# Diffusion Cocktail: Fused Generation from Diffusion Models

### v1

@@ 5 Conclusion
We propose Ditail, a training-free method to transfer content and style information between DMs. Ditail allows for novel generation by incorporating various DMs, giving users more control and flexibility and benefiting from the extensive collection of DMs. Ditail applies to style transfer tasks with the target style defined by a DM instead of images and has the promise of fine-grained manipulations such as prompt guidance strength and partial transfer for controllable image generations. Ditail is also generalizable to stylized image editing on real images. Our future work involves integrating Ditail to other emerging methods and architectures for better efficiency and controllability.
@@ Acknowledgments
This work is supported in part by the Shanghai Frontiers Science Center of Artificial Intelligence and Deep Learning at NYU Shanghai, NYU Shanghai Boost Fund, and NYU High Performance Computing resources and services.
Supplementary Material
6 Model Checkpoints Information
We manually select 5 DM checkpoints with distinctive style specialization from civitai  [^2] https://civitai.com. Fig. 11 lists the important information about these models, including the style keyword we assign and the tags provided by model uploaders. We also select some LoRA checkpoints from civitai and liblib  [^3] https://www.liblib.ai, whose information could be found in Fig. 12.
@@ 6 Model Checkpoints Information
7 Implementation Details
We implement the Ditail framework based on PyTorch and Diffusers  [^4] https://huggingface.co/docs/diffusers/index. Following prior works [37], we use Denoising Diffusion Implicit Models (DDIM) scheduler [34] with 1000 inversion steps and 50 sampling steps. Note that the inversion steps can be further reduced without a significant loss of output quality. In the project demo, we use 50 steps for both inversion and sampling and set $(\alpha,\beta)=(2.0,0.5)$ by default. With a Nvidia A10G-small GPU on HuggingFace, processing a single image takes roughly 15 seconds using 16-bit floating point precision.
8 Pseudocode for noise-pred-with-inj($\cdot$)
@@
2total_steps = 50
3thresh_res = 0.8 * total_steps
4thresh_attn = 0.5 * total_steps
@@
6
7# Input:
8#   - m: ResNet block to be forwarded
9#   - mask: Optional mask for local injection
10#   - x: Input hidden states
11# Output:
12#   - Forwarded hidden states
13def inj_forward_res(m, mask, x):
@@
18
19
20# Input:
21#   - m: Self-attention block to be forwarded
22#   - mask: Optional mask for local injection
23#   - x: Input hidden states
24# Output:
25#   - Forwarded hidden states
26def inj_forward_attn(m, mask, x):
@@
37
38
39# Input:
40#   - unet: U-Net model
41#   - x: Input noisy latents
42#   - t: Current time step
43#   - inj_ids: List of block IDs with injection
44# Output:
45#   - Denoised latent
46def noise_pred_with_inj(unet, x, t, inj_ids):
@@
51            if is_attn(m) and t > thresh_attn:
@@
Listing 1: Pseudocode for noise-pred-with-inj
9 Inversion Model Comparison
@@
One natural variant of Ditail style transfer is to invert a source image with non-original DMs. In this section, we perform both qualitative and quantitative investigations on the pros and cons of such a variant. As shown in Tab. 3, performing inversion with the target DM enhances structure preservation but degrades image fidelity and style. Some qualitative results are presented in Fig. 13. The first row demonstrates a failure case of inverting with source DM. This happens rarely when the source DM possesses a strong style, yet the target DM fails to denoise the strongly stylized latents properly and causes a collapse of content and structure. In such cases, inverting with target DMs can give better results. However, this also inevitably degrades the stylizing capability. As shown in the second and third rows, though inverting with the target DM ensures better structure preservation, this also leads to less stylized output. For real images, the quantitative results in Tab. 4 show similar patterns. Note that there is no source DM for real-world images, but its distribution can be best simulated by photo-realistic DMs, allowing more flexibility in the structure and thus resulting in more stylized output.
As an additional reference, we also attach the noisy latents inverted with different DMs in Fig. 14. We can observe that the noisy latents do exhibit some style-specific patterns. For example, the latents inverted with the Abstract DM are less colorful than the others. This is actually opposed to the patterns we observed from the output images, as inverting with the target DM does make the latents more “stylized”. Nevertheless, we hypothesize that target DMs tend to exhibit less capability of modifying the structure produced by other DMs, which may potentially lead to a better-preserved structure and a weakened style.
Inversion Model Scaling CLIP $\uparrow$ DINO $\downarrow$ $\text{FID}_{\text{tgt}}\downarrow$
Source $\alpha=1,\beta=0$ 0.1646 0.0425 196.41
$\alpha=2,\beta=0$ 0.1597 0.0354 194.02
$\alpha=4,\beta=0$ 0.1563 0.0267 209.23
$\alpha=8,\beta=0$ 0.1581 0.0245 228.05
Target $\alpha=1,\beta=0$ 0.1601 0.0288 204.87
$\alpha=2,\beta=0$ 0.1574 0.0262 202.09
$\alpha=4,\beta=0$ 0.1545 0.0236 212.77
$\alpha=8,\beta=0$ 0.1555 0.0232 230.18
Table 3: Quantitative comparison of using different DMs for inversion on generated images. The mean CLIP cosine similarity  [9, 26] on source images is 0.1589, and the mean FID [10] distance between source and target images is 227.27. The best-performed metric for each group is shown in bold.
Inversion Model Scaling CLIP $\uparrow$ DINO $\downarrow$ $\text{FID}_{\text{tgt}}\downarrow$
Realistic $\alpha=1$ 0.0761 0.0517 129.84
$\alpha=2$ 0.0772 0.0449 125.32
$\alpha=4$ 0.0882 0.0322 132.39
$\alpha=8$ 0.0950 0.0297 147.24
Target $\alpha=1$ 0.0750 0.0372 153.06
$\alpha=2$ 0.0781 0.0324 154.73
$\alpha=4$ 0.0872 0.0283 157.61
$\alpha=8$ 0.0952 0.0282 171.84
@@
Table 4: Quantitative comparison of using different DMs for inversion on real images. The mean CLIP cosine similarity  [9, 26] on source images is 0.0907, and the mean FID [10] distance between source and target images is 155.43. The best-performed metrics are shown in bold.
10 More Qualitative Results
We attach more qualitative results produced with the Ditail framework in this section. Fig. 15 presents the matrix view of cross-transfer across DMs of various styles (as in Fig. 5). Fig. 16 and Fig. 17 shows real image style transfer results with different model configurations and scaling factors. Fig. 18 ablates content injection and condition scaling on real images. Fig. 19 investigates the reversibility of Ditail style transfer. Fig. 20 explores the stability of the reconstructed content by performing Ditail multiple times. Please find the takeaways in the captions of each figure.
@@ 10 More Qualitative Results
Figure 13: Comparison of performing inversion with source and target DMs.
Figure 14: Visualizations of the noisy latents inverted by different DMs. We use the same source image as in Fig. 6.
@@
Figure 15: Demonstration of style transfer across various DMs. Similar to Fig. 5, images on the diagonal are those generated by the original DMs, whereas the others are the novel results. Ditail enables us to mix content and style information easily on images of all types.
@@
Figure 16: More style transfer results (DM only) on real images, which demonstrates Ditail’s high generalizability.
@@
Figure 17: More style transfer results (DM+LoRA) on real images with different scaling factors. During generation, trigger words for each LoRA are appended to the prompt, and we set the scale of LoRA weight to be 0.7, as recommended by LoRA uploaders. In most cases, we observe evident transformations in image style and content, while the structure of the source image is largely preserved. We also notice that the effect of the same LoRA on different checkpoints varies, which entails the DM-LoRA compatability problem.
Figure 18: Ablation studies on content injection and scaling factors for real images. We can see that applying either of them helps preserve the structure; otherwise, the output images will deviate from the original source content (i.e., close to vanilla text-to-image generation).
Figure 19: Results of transferring the novel images back to their original style. In general, the content deviation depends on 1) the content complexity and 2) the style gap between the source and target DMs. For example, the duckweed (complex) is subject to change after the two-step transfer; whereas the car (simple) remains largely the same. It is hard to restore the lighting as well (room example).
Figure 20: Results of reconstructing the input images with a single DM multiple times. In general, a larger $\alpha$ helps stabilize the content in the source image. However, it may lead to a weird coloring issue when the magnitude of the condition vector is too large.

### v2

@@ 6 Conclusion
We propose Ditail, a training-free method to transfer content and style information between multiple DMs. Ditail allows for image generation by incorporating various DMs, giving users more control and flexibility and benefiting from the extensive collection of DMs. Ditail applies to style transfer tasks with the target style defined by a DM instead of images. It promises fine-grained manipulations such as textual prompts and condition scaling for controllable image generations. Ditail is highly extensible to diverse application, such as novel-style image generation, prompt-based and collage-based image manipulation. Our future work involves integrating Ditail with other emerging methods and architectures for better efficiency and controllability.
@@ Acknowledgments
This work is supported by the Shanghai Frontiers Science Center of Artificial Intelligence and Deep Learning at NYU Shanghai, NYU Shanghai Boost Fund, and NYU High Performance Computing resources and services.
Appendix A Model Checkpoints Information
We manually select 5 DM checkpoints with distinctive style specialization from GEMRec-18K [8], which consists of 200 DM checkpoints from Civitai  [^2] https://civitai.com. Fig. 7 lists the important information about these models, including the style keyword and the tags provided by model uploaders. We also select some LoRA checkpoints from Civitai and Liblib  [^3] https://www.liblib.ai for demonstration, whose information could be found in Fig. 8.
@@ Appendix A Model Checkpoints Information
Appendix B Pseudocode for noise-pred-with-injection($\cdot$)
@@
2inference_steps = 50
3thresh_res = 0.8 * inference_steps
4thresh_attn = 0.5 * inference_steps
@@
6# Input:
7# - m: ResNet block to be forwarded
8# - mask: Optional mask for local injection
9# - x: Input hidden states
10# Output: Forwarded hidden states
11def inj_forward_res(m, mask, x):
@@
16
17# Input:
18# - m: Self-attention block to be forwarded
19# - mask: Optional mask for local injection
20# - x: Input hidden states
21# Output: Forwarded hidden states
22def inj_forward_attn(m, mask, x):
@@
33
34# Input:
35# - unet: U-Net model
36# - x: Input noisy latents
37# - t: Current time step
38# - inj_ids: List of block IDs with injection
39# Output: Denoised latent
40def noise_pred_with_inj(unet, x, t, inj_ids):
@@
45 if is_attn(m) and t >= thresh_attn:
@@
Appendix C Implementation Details
We implement Ditail using PyTorch and Diffusers  [^4] https://huggingface.co/docs/diffusers/index. We use Denoising Diffusion Implicit Models (DDIM) scheduler [36] with 50 inference steps for both inversion and sampling. We empirically set $(\alpha,\beta)=(3.0,0.5)$ by default to balance structure preservation and stylization performance. On a Nvidia A100 GPU, processing 300 images takes $\sim$25 minutes using FP16 precision (i.e., $\sim$5s/img).
Figure 9: Qualitative results for prompt-based content manipulation. We set the classifier-free guidance factor $\omega=7.5$ (left) and $15$ (right) for each image pair. The demonstration images are drawn from the LAION-5B [34] dataset with textual prompts generated by the captioner model from BLIP-Large [19]. Better view with color. Zoom in for best view.
Appendix D Prompt-based Image Manipulation
The qualitative results over various prompt types are shown in Fig. 9. We can observe that: 1) an empty prompt exhibits limited editing power compared to the others; 2) a larger CFG scale $\omega$ allows the target models to be more creative; 3) given that we enforce a strong structure emphasis through condition scaling and content injection, the editing prompts that change the number of objects usually give unsatisfactory results, whereas modifying the semantic class and/or colors of the subjects seems to be easier, even for completely unrelated prompts (e.g., man $\rightarrow$ chicken).
Appendix E Comparison of Injection Methods
Due to the memory burden of storing tensors (5.7 GB/img), we conduct evaluation over 300 randomly sampled (image, target style) pairs. The performance drops on CLIP, DINO, FID are 0.0013, 0.0018, 0.23, respectively, which are low enough to be negligible. Some qualitative results are shown in Fig. 10, and it can be seen that the differences are relatively minor, where the proposed injection method is more memory-friendly (4.7 MB/img) and efficient (i.e., it reconstructs the hidden states in parallel, so it only needs one noise sampling process).
Figure 10: Qualitative comparison of content injection methods. The middle column refers to content injection via cached tensors (i.e., $\{{\boldsymbol{h}}_{t}^{l}\},\{{\boldsymbol{q}}_{t}^{l}\},\{{\boldsymbol{k}}_{t}^{l}\}$). The right column refers to our proposed injection via reconstructing from latent ${\boldsymbol{z}}_{t}^{src}$.
Figure 11: Style mixing via checkpoints interpolation. Better view with color. Zoom in for best view.
Appendix F Qualitative Results for Style Interpolations
We present some qualitative results of interpolation between different styles in Fig. 11. By adjusting the interpolation factor $\gamma$, we can observe a smooth transition from one style to another. This technique allows us to tailor infinite target styles by merging existing DM checkpoints.
Appendix G Comparison of Different Inversion Models
@@
One natural variant of Ditail style transfer is to invert a source image with non-original DMs. We perform both qualitative and quantitative investigations on the pros and cons of the variants. In general, performing inversion with the target DM enhances structure preservation but degrades image fidelity and style. Some qualitative results are presented in Fig. 12. The first row demonstrates a rare failure case of inverting with source DM. This happens when the source DM possesses a strong style, yet the target DM fails to denoise the strongly stylized latents properly and leads to a collapse of content and structure. In such cases, inverting with target DMs can give better results. However, this also inevitably degrades the stylizing capability. As shown in the second and third rows, inverting with the target DM preserves the content, but also results in less stylized output.
[…]
Photo-realistic $\alpha=1,\beta=1$ 0.2499 0.0368 127.64
$\alpha=2,\beta=1$ 0.2466 0.0432 120.65
$\alpha=4,\beta=1$ 0.2499 0.0385 124.60
$\alpha=8,\beta=1$ 0.2505 0.0334 146.34
Target Style $\alpha=1,\beta=1$ 0.2523 0.0332 130.80
$\alpha=2,\beta=1$ 0.2518 0.0380 124.34
$\alpha=4,\beta=1$ 0.2525 0.0339 128.36
$\alpha=8,\beta=1$ 0.2516 0.0312 149.57
@@
Table 2: Quantitative comparison of inversion with different DMs on real images. The mean CLIP cosine similarity [10, 28] on source images is 0.2533, and the FID [11] between source and target images is 154.89. The best-performed metrics are shown in bold.
@@
Figure 12: Comparison of performing inversion with source and target DMs. Best view with color.
Figure 13: Qualitative comparison of performing inversion with photo-realistic and generic DMs. Zoom in for best view.
For real images, the quantitative results in Tab. 2 demonstrates the trade-off between structure and style. We can see that performing inversion with the target DM tend to enhance structure preservation but degrades the stylization performance. Given that there is no source DM for real-world images, its image distribution can be better approximated by photo-realistic DMs, allowing more flexibility in the structure and thus resulting in better stylized output. We also compare the qualitative results of using photo-realistic (Realistic Vision) or generic (Stable Diffusion 1.5) DMs in Fig. 13 for inversion, where the differences are not significant. By default, we use photo-realistic DMs for real-world images as it is further fine-tuned on this domain.
@@
Figure 14: Demonstration of novel image generation across various DMs. Images on the diagonal are generated by the original DMs, whereas the others are the novel results. Ditail enables us to mix content and style information easily on images of all types.
Appendix H More Qualitative Results for Style Transfer
We attach more qualitative results produced with Ditail. Fig. 14 presents the matrix view of cross-transfer across DMs of various styles. Fig. 15 and Fig. 16 shows real image style transfer results with different model configurations and scaling factors. By taking advantage of the vast collection of DM/LoRA checkpoints, Ditail is able to transfer and mix content/style information in a training-free manner. The presented qualitative results demonstrate its high applicability and generalization ability. In theory, Ditail can generate any subjects of any style, as long as we have the corresponding source images and target DM checkpoints.
@@ Appendix H More Qualitative Results for Style Transfer
Figure 15: More style transfer results (DM only) on real images. Better view with color.
@@
Figure 16: More style transfer results (DM+LoRA) on real images with different scaling factors. During generation, trigger words for each LoRA are appended to the prompt, and we set the scale of LoRA weight to be 0.7, as recommended by LoRA uploaders. We can observe evident transformations in image style and content, while the structure of the source image is largely preserved. We also notice that the effect of the same LoRA on different checkpoints varies, which entails the DM-LoRA compatability problem. Better view with color.